# Universidad Andres Bello, Facultad de Ciencias Exactas, Departamento de Ciencias Físicas - Instituto de Astrofísica, Autopista Concepcion-Talcahuano 7100, Talcahuano, Chile Instituto de Física, Universidad de Antioquia, Calle 70 52-21, Medellín, Colombia

### v1

@@ Abstract
Context. The Bulge Globular Cluster NGC 6316 has been the subject of a limited amount of previous chemical studies besides metallicity. As part of the bulge Cluster APOgee Survey (CAPOS), we can now improve our knowledge of the chemistry and nature of this cluster. CAPOS makes use of high-resolution, high Signal-to-Noise Ratio Near-Infrared spectroscopy, which is capable of penetrating the substantial dust towards the Galactic Bulge - a significant obstacle for this cluster due to its high reddening.
Aims. We aim to conduct the most robust chemical study to date for NGC 6316. Using high-resolution, high Signal-to-Noise Ratio spectra of 10 RGB members, we aim to derive abundances for a number of elements with a variety of nucleosynthetic origins, most of which have never been studied before in this cluster.
@@
Results. We determined, for the first time, high resolution spectroscopic abundances for C, N, O, Mg, Al, Si, P, K, Ca, Ti, V, Cr, Mn, Fe, Ni and Ce for this cluster. We obtained a mean metallicity [Fe/H] = 0.87 0.02, finding no indication of an intrinsic metallicity spread. Our metallicity agrees with the most recent values from other studies, revising earlier values that were 0.5 dex metal-richer. With this new value, this cluster, long believed to be a member of the classical metal-rich group of bulge GCs around -0.5, now falls in the dominant bulge globular cluster peak around [Fe/H] = -1. The cluster presents a clear C-N anticorrelation. We also found a [/Fe] = 0.3 0.02. Our abundances show similar behaviour to other in situ globular clusters with comparable metallicity. We obtained and by isochrone fitting, in good agreement with the recent determinations from other works. We derive an overall metallicity by isochrone fitting, in agreement with our abundance determination. According to the mean [Mg/Fe] and [Al/Fe] abundances from first population stars, NGC 6316 is an in-situ globular cluster, in accordance with various dynamical classifications.
@@ Conclusions.
Globular Clusters (GCs) are among the oldest objects in the universe and also present Multiple Populations (MPs). These characteristics make GCs very interesting objects to study in the context of galactic and cluster formation. While studies related to the age of GCs have existed for quite some time now (Sandage, 1953), studies of MPs are relatively new. This concept, described as significant intrinsic abundance variations within a GC, emerged some 15 years ago when Carretta et al. (2009) suggested that this phenomenon must be fundamentally related to the mechanisms of formation and early evolution of GCs since, at that time, all the GCs studied (e.g., Bedin et al. (2004), among others) presented MPs. To this day, many GCs have been studied by combining different photometric (Frelijj et al., 2017, 2022; Piotto et al., 2015; Milone et al., 2017) and spectroscopic (Carretta et al., 2009; Mészáros et al., 2020) techniques. Great advances have been made, such as the creation of large databases (e.g., Baumgardt111https://people.smp.uq.edu.au/HolgerBaumgardt/globular/ and Harris, 1996 Galactic GCs databases), identification of chemical correlations and anti-correlations, differentiation of the populations through them (Carretta et al. (2010), Mészáros et al. (2020)), and even the identification of a new type of GC, Type II (Milone et al., 2017). However, the lack of a viable formation scenario for GCs and MPs, e.g. to account for the mass budget problem, as well as an explanation for the concomitant existence of at least some true Simple Stellar Population GCs (E3: Salinas, & Strader (2015); Monaco et al. (2018), Ruprecht 106: Villanova et al. (2013); Dotter et al. (2018); Frelijj et al. (2021), Ter7 & Pal12: Cohen (2004); Tautvaišienė et al. (2004); Sbordone et al. (2007)), among other issues, only proves that there is still a long way to go to fully understand the formation mechanisms of GCs. The only way to solve these questions is to obtain both more data, to help constrain formation scenarios, as well as additional theoretical ideas to test.
@@
The order of the columns is as follows: Star ID, Right Ascension(J2000), Declination(J2000), GAIA magnitude in G, GBP and GRP, 2MASS magnitude in J, H and Ks, Absolute Proper Motion in RA and DEC from GAIA, Heliocentric RV from ASPCAP, and SNR. IDs marked with an * indicate spectra not obtained by CAPOS.
@@
Given that many of these questions remain unanswered after studying a large number of primarily halo GCs, a promising direction for future studies is to turn to the study of GCs in other Galactic components. The Galaxy possesses no less than three GC systems: the halo, bulge and disk systems (e.g. Massari, Koppelman, & Helmi, 2019). Indeed, these latter two contain fully 2/3 of the total number of Galactic GCs (Belokurov & Kravtsov, 2024). The study of bulge globular clusters (BGCs) in particular opens up the regime of high metallicity to MP studies since they include the highest metallicity of all GCs, and thus help verify metallicity trends so far only hinted at (e.g. Muñoz et al., 2020). Nevertheless, most of the Bulge/Disk (B/D) GC systems have not been studied in the detail required to quantify or even explore their MP nature, which demands either deep UV photometry with high spatial resolution, exemplified by the HST UV Legacy Survey of GCs Piotto et al. (2015), or high resolution high Signal-to-Noise Ratio (SNR) spectroscopy of large samples. Unfortunately, due to the vast amount of extinction in the plane and towards the center of the Galaxy (e.g Geisler et al., 2021), UV observations of the B/D GC systems are severely limited, so that the HST UV Legacy Survey includes only a very small number of these objects. Optical spectroscopy is similarly limited. However, the high reddening problem that previously prevented us from studying the B/D has been quite successfully addressed with the help of near-infrared (NIR) instruments. Surveys like the VISTA Variables in the Via Lactea (VVV) (Minniti et al., 2010) with its eXtension (VVVX)(Saito et al., 2024), and the Apache Point Observatory Galactic Evolution Experiment of the Sloan Digital Sky Survey III (SDSS-III APOGEE) (Majewski et al., 2017) with its successor (SDSS-IV APOGEE-2) capitalize on this capability to great success. However, despite these large surveys, the VVV survey lacks UV observations while the SDSS-IV survey did not prioritize B/D GCs. Mészáros et al. (2020) present the SDSS-IV sample of 44 GCs, of which only 2 are bonafide BGCs according to Massari, Koppelman, & Helmi (2019) and have a large enough sample of well-observed members and low enough reddening to qualify for their study. This is ¡4% of the total number of BGCs known, a disconcertingly low value.
@@
In order to take full advantage of the wealth of astrophysical detail these key objects can provide, as complete a sample as possible is essential. To this end, ”the bulge Cluster APOgee Survey” (CAPOS) (Geisler et al., 2021) was created. This survey was designed to obtain detailed abundances and kinematics for as complete a sample of bona fide BGCs as possible using the unique advantages of APOGEE in order to exploit their extraordinary Galactic archaeology attributes. In addition to providing reliable multi-element abundances for a number of B/D GCs (many for the first time), this ongoing project has published a series of important results: e.g. confirmation of the scenario that Main Bulge and Main Disk GCs, formed in situ, have [Si/Fe] abundances slightly higher than their accreted counterparts at the same metallicity (Geisler et al., 2021, CAPOS I Paper); the existence of the characteristic N-C anticorrelation, and Al-N correlation for the first time in FSR 1758 (Romero-Colmenares et al., 2021, CAPOS II Paper); a Ce-N correlation in Tonantzintla 2 (Fernández-Trincado et al., 2022, CAPOS III Paper); and a N-C anticorrelation for the first time in NGC 6558 (González-Díaz et al., 2023, CAPOS IV Paper). These studies, which are providing valuable information about B/D GCs, demonstrate that it is possible to obtain and analyze excellent data that, until very recently, was very difficult to obtain.
@@
Although there is already an APOGEE pipeline that derives chemical abundances and other parameters like radial velocity for BGCs, namely the APOGEE Stellar Parameters and Chemical Abundance Pipeline (ASPCAP, García Pérez et al. (2016)), it is known that the pipeline does not perform optimally for stars with extreme types of “non-standard” elemental abundance patterns, namely so-called Second Population (2P) GC stars - those that have been enriched by some as yet-to-be-determined process (Holtzman et al., 2015; Jönsson et al., 2018). Fernández-Trincado et al. (2020) found significant differences in certain abundance ratios compared to ASPCAP, in the range of dex higher than ASPCAP values for [N, O, Al, Si, Ce/Fe] to dex lower for [O, Mg/Fe] in 2P stars. CAPOS I showed that even the ASPCAP Fe abundances of 2P stars deviates from that of their first population (1P) cousins, with the former offset to somewhat higher metallicity values within a given GC. Hence, different abundance determination techniques are necessary for 2P stars at least, and preferably one should also perform the same analysis on 1P stars in order to ensure homogeneity.
@@
NGC 6316 is regarded as one of the classic metal-rich BGCs, with a metallicity of 0.45 listed by Harris (1996, 2010 update). However, due to the extinction problem described above, no previous high resolution spectroscopic study has been performed on it, and it is not part of the HST UV Legacy survey, so nothing is known of its MP nature. Its metallicity has only been estimated via relatively uncertain photometric or low resolution spectroscopic means (e.g. Conroy et al. 2018, [Fe/H] = -0.87). The best of these, until now, is the Calcium triplet study of Vásquez et al. (2018), who found a mean metallicity of 0.37 0.17 from only 3 stars, in good agreement with the Harris catalog. Most recently, Deras et al. (2023) used deep HST photometry to estimate the age and metallicity, deriving [Fe/H] 0.9 and 13.1 0.5 Gyr, making it ”another extremely old relic of the assembly history of the Galaxy” according to them, but they lament the lack of a high-quality metallicity determination for this GC, which of course is critical for deriving an accurate age. Note that their value is about 0.5 dex lower than previous estimates, further demanding a more robust metallicity derivation for this cluster.
According to the Baumgardt database (and references therein), NGC 6316 lies at J2000 coordinates RA: 259.15542 deg and DEC: 28.14011 deg, it is located at a heliocentric distance of kpc and Galactocentric distance of 3.16 0.36 kpc, has mean absolute proper motions (PMs) of pmRA = 4.974 0.028 masyr-1 and pmDEC = 4.605 0.027 masyr-1, a mean heliocentric radial velocity (RV) of kmsec-1, a mass of M⊙, and a tidal radius (TR) pc. Deras et al. (2023), provides a foreground reddening of E(B-V) .
Massari, Koppelman, & Helmi (2019) classify it as a Main Bulge GC, Pérez-Villegas et al. (2020) as Thick Disk, Callingham et al. (2022) as most likely originating from Kraken, and both Belokurov & Kravtsov (2024) and Chen & Gnedin (2024) as in situ. Geisler et al. (in prep) give NGC 6316 an 81% probability of being an in situ GC and a 51% probability of being a BGC as opposed to Disk GC, and finally classify it as a bonafide BGC.
@@
The final sample is depicted in all plots as bigger circles with a color code depending on their SNR.
@@ 2 Data and target selection
The data consists of high-resolution spectra (R 22500), centered in the near-infrared (H-Band), ranging from 1.51 m to 1.70 m observed through the APOGEE spectrograph (Wilson et al., 2019) on the 2.5m Irénée du Pont telescope at Las Campanas Observatory (Bowen & Vaughan, 1973), used by the APOGEE survey to observe the Southern Hemisphere (APOGEE-2S, Santana et al., 2021).
NGC 6316 was observed as part of the CAPOS survey (Geisler et al., 2021) through the Contributed APOGEE-2S CNTAC program CN2019B. Since the observations were obtained prior to the release of Gaia DR2 data, targets were based on existing position, color-magnitude and radial velocity information.
The final cluster members were selected from the final version of the APOGEE-2 catalog published in December 2021 as part of the Data Release 17 of the Sloan Digital Sky Survey (SDSS) (Abdurro’uf et al., 2022), which is available in the SDSS IV Science Archive Server222https://www.sdss4.org/dr17/irspec/spectro_data/.
Based on the information given in this catalog, we selected targets as potential members of NGC 6316 through 5 criteria:
@@
- We took the mean PMs values from GAIA (Gaia Collaboration et al., 2016, 2021), attached to the APOGEE catalog, and selected all the stars inside a radius of 0.5 masyr-1 from the mean value given by Baumgardt database. (Blue dots in the top right panel of Figure 1)
- Taking into account the mean RV and [Fe/H] of the cluster given by Baumgardt database and Deras et al. (2023), we selected a range of stars whose RVs, determined by the APOGEE survey, were around this value, ranging from kmseg-1, and whose metallicities [Fe/H] from ASPCAP were close to the value given by Deras et al. (2023) for the cluster. (Green dots and other non blue circles in the bottom left panel of Figure 1)
- A Color-magnitude diagram with GAIA photometry in vs was made to ensure that all the potential targets lay along the main Red Giant Branch (RGB). (Colored points in bottom right panel of Figure 1)
@@
A cross-match of our 10 targets with the Vasiliev & Baumgardt (2021) catalog of NGC 6316 determined that, according to the criteria used in their study, each star has a probability of over 80% of being part of the cluster, adding further reliability that our membership selection is very robust.
Table 1 showlists the member parameters. These selections led to the previously mentioned total of 10 targets, of which 8 were observed as part of the CAPOS survey (Geisler et al., 2021), and the remaining 2 were observed under the main APOGEE survey (Majewski et al., 2017). The spectra of these 10 targets were downloaded from the SDSS IV Science Archive Server444https://data.sdss.org/sas/dr17/apogee/spectro/aspcap/dr17/synspec_rev1/. All the obtained spectra had already undergone reduction using the APOGEE-2 data reduction pipeline as described in Nidever et al. (2015).

### v2

# CAPOS: The bulge Cluster APOgee Survey VII. First detailed chemical analysis of NGC 6316

@@ Abstract
Context. The Bulge Globular Cluster NGC 6316 has had few previous chemical studies beyond metallicity. As part of the bulge Cluster APOgee Survey (CAPOS), we can now improve our knowledge of the chemistry and nature of this cluster. CAPOS makes use of high-resolution, high signal-to-noise ratio near-infrared spectroscopy, which is capable of penetrating the substantial dust towards the Galactic Bulge - a significant optical obstacle for this cluster due to its high reddening.
Aims. We aim to conduct the most robust chemical study to date for NGC 6316 by deriving abundances for a number of elements with a variety of nucleosynthetic origins, most of which have never been studied before in this cluster.
@@
Results. We determined, for the first time, high-resolution spectroscopic abundances for C, N, O, Mg, Al, Si, P, K, Ca, Ti, V, Cr, Mn, Fe, Ni and Ce for this cluster. We obtained a mean , finding no indication of an intrinsic metallicity spread. Our metallicity agrees with the most recent values from other studies, revising earlier values that were 0.5 dex metal-richer. With this new value, this cluster, long believed to be a member of the classical metal-rich group of bulge GCs around -0.5, now falls in the dominant bulge globular cluster peak around . The cluster presents a clear C-N anti-correlation and [/Fe]=0.310.02. Our abundances show similar behaviour to other in-situ globular clusters with comparable metallicity. An isochrone fitting gave us , a higher value than any other from the literature for this cluster since we also estimated , in good agreement with determinations from other works; . We derive an overall metallicity , in agreement with our abundance determination.
@@ Conclusions.
Globular Clusters (GCs) are among the oldest objects in the universe. This characteristic make GCs very interesting objects to study in the context of galaxy and cluster formation. Because of this, they have been the main target of a vast amount of research, which has yielded unprecedented results in these fields. These include cluster distances, reddening, and ages (Dotter et al., 2010; Cohen & Sarajedini, 2012; VandenBerg et al., 2013), luminosity and mass functions (Paust et al., 2010), structural parameters and mass segregation measurements (Goldsbury, Heyl, & Richer, 2013), mass loss (Caloi & D’Antona, 2008; Momany et al., 2012) and studies in multiple populations (MPs; Carretta et al., 2009a; Villanova et al., 2013; Piotto et al., 2015; Milone et al., 2017), among others. Significant advances have been made in our understanding of GCs and the formation history of the Milky Way. However, a consistent theoretical framework is still lacking, and key questions remain regarding the fundamental nature of GCs and the broader history of our Galaxy.
@@
The order of the columns is as follows: Star ID, Right Ascension(J2000), Declination(J2000), GAIA DR3 (Gaia Collaboration et al., 2016, 2021) magnitude in G, GBP and GRP, 2MASS magnitude in J, H and Ks, Absolute Proper Motion in RA and DEC from GAIA DR3, Heliocentric RV from ASPCAP, and SNR. IDs marked with an * indicate spectra not obtained by CAPOS.
@@
Given that many questions remain unanswered after studying a large number of primarily halo GCs, a promising direction to focus future studies is to turn to the study of GCs in other Galactic components. The Galaxy possesses no less than three GC systems: the halo, bulge and disk systems (e.g. Massari, Koppelman, & Helmi, 2019). Indeed, these latter two contain fully 2/3 of the total number of Galactic GCs (Belokurov & Kravtsov, 2024). The study of bulge globular clusters (BGCs) in particular opens up the regime of high metallicity to MP studies since they include the highest metallicity of all GCs, and thus help verify metallicity trends so far only hinted at (e.g. Muñoz et al., 2020). Nevertheless, most of the Bulge/Disk (B/D) GC systems have not been studied in the detail required to quantify or even explore their MP nature, which demands either deep UV photometry with high spatial resolution, or high resolution high Signal-to-Noise Ratio (SNR) spectroscopy of large samples. Unfortunately, due to the vast amount of extinction in the plane and towards the centre of the Galaxy (e.g. Gonzalez et al., 2012), UV observations of the B/D GC systems are severely limited, so that the HST UV Legacy Survey (Piotto et al., 2015) includes only a very small number of these objects. Optical spectroscopy is similarly limited. However, the high reddening problem that previously prevented us from studying the B/D can be quite successfully addressed with the help of near-infrared (NIR) instruments. Surveys like the VISTA Variables in the Via Lactea (VVV) (Minniti et al., 2010) with its eXtension (VVVX)(Saito et al., 2024), and the Apache Point Observatory Galactic Evolution Experiment of the Sloan Digital Sky Survey III (SDSS-III APOGEE) (Majewski et al., 2017) with its successor (SDSS-IV APOGEE-2) capitalize on this capability to great success. However, despite these large surveys, the VVV survey lacks UV observations while the SDSS-IV survey did not prioritize B/D GCs. Mészáros et al. (2020) present the SDSS-IV sample of 44 GCs, of which only 2 are bona fide BGCs according to Massari, Koppelman, & Helmi (2019) and have a large enough sample of well-observed members and low enough reddening to qualify for their study. This is ¡4% of the total number of BGCs known, a disconcertingly low value.
@@
Despite the existence of multiple individual studies of BGCs (e.g. Origlia et al., 2005; Valenti, Origlia, & Rich, 2011; Villanova et al., 2017; Muñoz et al., 2017, among others), a homogeneous study, involving a large number of BGCs is still lacking. In order to take full advantage of the wealth of astrophysical detail these key objects can provide, as complete a sample as possible is essential. To this end, ”the bulge Cluster APOgee Survey” (CAPOS) (Geisler et al., 2021) was created. This survey was designed to obtain detailed abundances and kinematics for as complete a sample of bona fide BGCs as possible using the unique advantages of APOGEE in order to exploit their extraordinary Galactic archaeology attributes. In addition to providing reliable multi-element abundances for a number of B/D GCs (many for the first time), this ongoing project has published a series of important results: e.g. confirmation of the scenario that Main Bulge and Main Disk GCs, formed in situ, have [Si/Fe] abundances slightly higher than their accreted counterparts at the same metallicity (Geisler et al., 2021, CAPOS I Paper); the existence of the characteristic N-C anti-correlation, and Al-N correlation for the first time in FSR 1758 (Romero-Colmenares et al., 2021, CAPOS II Paper); a Ce-N correlation in Tonantzintla 2 (Fernández-Trincado et al., 2022, CAPOS III Paper); a N-C anti-correlation for the first time in NGC 6558 (González-Díaz et al., 2023, CAPOS IV Paper), and detailed studies of HP 1 (Henao et al., 2025, CAPOS V Paper) and NGC 6569 (Barrera et al., 2025, CAPOS VI Paper). These studies, which are providing valuable information about B/D GCs, follow a coordinated methodology that provides homogeneous studies, allowing a uniform analysis of the data and results.
@@
Although there is already an APOGEE pipeline that derives chemical abundances and other parameters like radial velocity for BGCs, namely the APOGEE Stellar Parameters and Chemical Abundance Pipeline (ASPCAP, García Pérez et al. (2016)), it is known that the pipeline does not perform optimally for stars with extreme types of “non-standard” elemental abundance patterns, namely so-called Second Population (2P) GC stars - those that have been enriched by some as yet-to-be-determined process (Holtzman et al., 2015; Jönsson et al., 2018). Fernández-Trincado et al. (2020) found significant differences in certain abundance ratios compared to ASPCAP, in the range of dex higher than ASPCAP values for [N, O, Al, Si, Ce/Fe] to dex lower for [O, Mg/Fe] in 2P stars. CAPOS I showed that even the ASPCAP Fe abundances of 2P stars deviates from that of their first population (1P) cousins, with the former offset to somewhat higher metallicity values within a given GC. This was strongly corroborated by Geisler et al. (2025) recent analysis of APOGEE results for all CAPOS GCs. Hence, different abundance determination techniques are necessary for 2P stars at least, and preferably one should also perform the same analysis on 1P stars in order to ensure homogeneity.
@@
NGC 6316 is regarded as one of the classic metal-rich BGCs, with a metallicity of 0.45 listed by Harris (1996, 2010 update, hereafter H10). However, due to the extinction problem described above, no previous high resolution spectroscopic study has been performed on it, and it is not part of the HST UV Legacy survey, so nothing is known of its MP nature. Its metallicity has only been estimated via relatively uncertain photometric or low resolution spectroscopic means, e.g. Zinn & West (1984, hereafter ZW84)(Pseudo-equivalent width transformed to : ), Armandroff & Zinn (1988, hereafter AZ88)(Calcium triplet: ; averaged with ZW84: ), Valenti, Ferraro, & Origlia (2007, hereafter VFO07)(mid-resolution photometry: ), Carretta et al. (2009c, hereafter C09)(UVES recalibration of the 1984 : ), Conroy et al. (2018)(stellar population models: ). The best spectroscopic study, until now, is the Calcium triplet study of Vásquez et al. (2018), who found a mean metallicity of , and using [Fe/H] calibrations from Dias et al. (2016), Saviane et al. (2012) and Vásquez et al. (2015), respectively, from only 3 stars, in good agreement with the H10 catalogue. Most recently, Deras et al. (2023, hereafter D23) used deep HST photometry to estimate the age and metallicity, deriving [Fe/H] 0.9 and 13.1 0.5 Gyr, making it ”another extremely old relic of the assembly history of the Galaxy” according to them, but they lament the lack of a high-quality metallicity determination for this GC, which of course is critical for deriving an accurate age. Note that their value is about 0.5 dex lower than previous estimates, further demanding a more robust metallicity derivation for this cluster.
According to the Baumgardt database111https://people.smp.uq.edu.au/HolgerBaumgardt/globular/ (and references therein), NGC 6316 lies at J2000 coordinates RA: 259.15542 deg and DEC: 28.14011 deg, it is located at a heliocentric distance of kpc and Galactocentric distance of 3.16 0.36 kpc, has mean absolute proper motions (PMs) of pmRA = 4.974 0.028 masyr-1 and pmDEC = 4.605 0.027 masyr-1, a mean heliocentric radial velocity (RV) of kmsec-1, a mass of M⊙, and a tidal radius (TR) pc. D23 provides a foreground reddening of E(B-V) .
Massari, Koppelman, & Helmi (2019) classify it as a Main Bulge GC, Pérez-Villegas et al. (2020) as Thick Disk, Callingham et al. (2022) as most likely originating from Kraken, and both Belokurov & Kravtsov (2024) and Chen & Gnedin (2024) as in situ. Geisler et al. (2025) give NGC 6316 an 81% probability of being an in situ GC and a 51% probability of being a BGC as opposed to Disk GC, and finally classify it as a bona fide BGC.
@@
The final sample is depicted in all plots as bigger circles with a colour code depending on their SNR.
@@ 2 Data and target selection
The data consists of high-resolution spectra (R 22500), centred in the near-infrared (H-Band), ranging from 1.51 m to 1.70 m observed through the APOGEE spectrograph (Wilson et al., 2019) on the 2.5m Irénée du Pont telescope at Las Campanas Observatory (Bowen & Vaughan, 1973), used by the APOGEE survey to observe the Southern Hemisphere (APOGEE-2S, Santana et al., 2021).
NGC 6316 was observed as part of the CAPOS survey (Geisler et al., 2021) through the Contributed APOGEE-2S CNTAC program CN2019B. Since the observations were obtained prior to the release of Gaia DR2 data, targets were based on existing position, colour-magnitude and radial velocity information.
The final cluster members were selected from the final version of the APOGEE-2 catalogue published in December 2021 as part of the Data Release 17 of the Sloan Digital Sky Survey (SDSS) (Abdurro’uf et al., 2022), which is available in the SDSS IV Science Archive Server222https://www.sdss4.org/dr17/irspec/spectro_data/.
Based on the information given in this catalogue, we selected observed targets as potential members of NGC 6316 through 5 criteria:
@@
- We took the mean PMs values from GAIA DR3, attached to the APOGEE catalogue, and selected all the stars inside a radius of 0.5 masyr-1 from the mean value given by Baumgardt database. (Blue dots in the top right panel of Figure 1)
- Taking into account the mean RV and [Fe/H] of the cluster given by Baumgardt database and D23, we selected a range of stars whose RVs, determined by the APOGEE survey, were around this value, ranging from kmsec-1, and whose metallicities [Fe/H] from ASPCAP were close to the value given by D23 for the cluster. (Green dots and other non blue circles in the bottom left panel of Figure 1)
- A colour-magnitude diagram with GAIA DR3 photometry in vs was made to ensure that all the potential targets lay along the main Red Giant Branch (RGB). (coloured points in bottom right panel of Figure 1)
@@
A cross-match of our 10 targets with the Vasiliev & Baumgardt (2021) catalogue of NGC 6316 determined that, according to the criteria used in their study, each star has a probability of over 80% of being part of the cluster, adding further reliability that our membership selection is very robust.
Table 1 lists the member parameters. These selections led to the previously mentioned total of 10 targets, of which 8 were observed as part of the CAPOS survey (Geisler et al., 2021), and the remaining 2 were observed under the main APOGEE survey (Majewski et al., 2017). The spectra of these 10 targets were downloaded from the SDSS IV Science Archive Server444https://data.sdss.org/sas/dr17/apogee/spectro/aspcap/dr17/synspec_rev1/. All the obtained spectra had already undergone reduction using the APOGEE-2 data reduction pipeline as described in Nidever et al. (2015).